### v1

@@ 4.2 Improving accuracy of MLP via Fine-tuning
Fine tuning was carried out on our curated dataset using the multi-head approach as implemented in the MACE code starting with version 0.3.7. We used MACE-MP-0b as one head and our PBE+D3(BJ) [62, 63, 65, 64] calculations as the other head. The ratio of weights for the loss functions for energy, forces and stresses was 1:10:100 and we tuned the MLP training for 2500 epochs. It is important to note that during the training process, MACE employs a FPS approach to select configurations from the original MPtrj training set [35] and adds them to our data. Table 3 shows the final convergence results for the fine tuning for the two models trained, MACE-MP-MOF0 and MACE-MP-MOF0-v2. Due to the relatively large number of atoms in our training set per MOF, training was carried out on 16 A100@40GiB RAM NVIDIA GPUs.

### v3

@@ 4.2 Improving accuracy of MLP via Fine-tuning
Fine tuning was carried out on our curated dataset using the multi-head approach as implemented in the MACE code starting with version 0.3.7. We used MACE-MP-0b as one head and our PBE+D3(BJ) [67, 68, 70, 69] calculations as the other head. The ratio of weights for the loss functions for energy, forces and stresses was 1:10:100 and we tuned the MLP training for 2500 epochs. It is important to note that during the training process, MACE employs a FPS approach to select configurations from the original MPtrj training set [36] and adds them to our data. Table 3 shows the final convergence results for the fine tuning for the two models trained, MACE-MP-MOF0 and MACE-MP-MOF0-v2. Due to the relatively large number of atoms in our training set per MOF, training was carried out on 16 A100@40GiB RAM NVIDIA GPUs. The hyperparameters for fine tuning the model, are inherited from the original MACE-MP-0b model (L=1, $r\_max=6.0$, hidden irreps 128x0e+128x1o, 2 layers, each with correlation order: 3 (body order: 4) and spherical harmonics up to: l=3, see [34] for the full list), Specifically hyperparameters for fine tuning are: learning rate 0.005, ema decay 0.995, batch size 8.